Isopropyl alcohol,CH3CHOHCH3, is used in rubbing alcohol mixtures. How much heat is absorbed by alcohol if 10.0grams evaporates. The heat of vaporation is 42.1 kJ/mol
1
Expert's answer
2020-02-03T07:45:57-0500

n(CH3CHOHCH3) = m(CH3CHOHCH3)/m(CH3CHOHCH3) = 10.0g/60g/mol = 0.1667mol.

Q = L(CH3CHOHCH3) * n(CH3CHOHCH3) = 42.1 kJ/mol * 0.1667mol = 7.02kJ.